[PIPE2D-1132] b1 detectormap has incorrect wavelength solution Created: 15/Dec/22  Updated: 16/Dec/22  Resolved: 16/Dec/22

Status: Done
Project: DRP 2-D Pipeline
Component/s: None
Affects Version/s: None
Fix Version/s: None

Type: Bug Priority: Normal
Reporter: hassan Assignee: hassan
Resolution: Done Votes: 0
Labels: EngRun, calib-mgmt
Remaining Estimate: Not Specified
Time Spent: Not Specified
Original Estimate: Not Specified

Attachments: File bootstrap-20221214.md     File calib-generation-20221214-notes.md     PNG File NaI_doublet_83259.png     PNG File O5577_83259_b1b3_fixed.png     PNG File O5577_83259_b1b3.png     PNG File O8347_r1r3.png    
Issue Links:
Relates
relates to PIPE2D-1133 Correct fiber throughputs Done
Sprint: PreEngRun09Dec
Reviewers: price

 Description   

Fix problem reported with the current b1 detectormap on slack: https://sumire-pfs.slack.com/archives/C3AU3VCMU/p1670340234071919?thread_ts=1669909879.952659&cid=C3AU3VCMU in which the wavelength solution for O 5577 line differs significantly in b1 compared with b3. See also O5577_83259_b1b3.png .

In addition, now that we have simultaneous exposures of arc and scienceTrace data with SM1 and SM3, regenerate fiberProfiles and detectormaps in b and r using visits with such exposures.



 Comments   
Comment by hassan [ 15/Dec/22 ]

Regenerated b1, r1, m1, b3, r3 fiberProfiles and detectormaps, and the m3 detectormap, using better bootstrap solutions. Bootstrap detectormaps (and build notes} are now stored under the detectorMap/bootstrap subdirectory of the drp_pfs_data repo as a basis when building future detectormaps.

Reprocessed visit 83259 using update calibs. Wavelength solution for b1 much improved. See O5577_83259_b1b3_fixed.png. Other lines across b and r also look consistent across SM1 and SM3 (see NaI_doublet_83259.png and O8347_r1r3.png ).

New fiberProfiles and detectormaps installed in Hilo in new official calib location of /work/drp/CALIB (note: no longer a symlink).

Calib build notes attached and also included in /work/drp/CALIB/notes directory.

Comment by price [ 16/Dec/22 ]

We also need the calibs installed on the Tiger cluster.

Thanks for the notes, which should be very handy going forward.

The quartz spectra in the fiberProfiles have 3x higher flux on one side of the PFI than the other, so we need to fix that or it will adversely affect sky subtraction and flux calibration. If there isn't another set with more suitable illumination, I'm hoping we can manually scale them to match the actual throughput from the sky (via a twilight exposure, maybe?).

Comment by rhl [ 16/Dec/22 ]

We have to sort out the fibre throughputs.  Twilight sky is one option (with field rotation!) but it's a bit tricky to get.  Another option is to use the bright sky lines (e.g. 5577 in the blue, and reasonably-isolated sets of Meinel lines in the red).  They should be uniform across the field of view to a percent or two, and we can beat that down as sqrt(N).  It'd be good to confirm that both approaches give the same result.

Comment by hassan [ 16/Dec/22 ]

Thanks.

I'll close this ticket based on the CALIB created, and file a new ticket to improve the fiber throughputs.

Comment by hassan [ 16/Dec/22 ]

Latest calib update also copied to Tiger: /projects/HSC/PFS/Subaru/CALIB (no symlink).

Generated at Sat Feb 10 16:03:12 JST 2024 using Jira 8.3.4#803005-sha1:1f96e09b3c60279a408a2ae47be3c745f571388b.